f the volume of ocean water is 4.18*10^20L contains 4.3 percent sodium chloride ("salt") by mass and the density of the water is 1.56 g/mL, calculate the total mass of sodium chloride. (1 ton = 2000 lb, 1 lb = 453.6 g)
1
Expert's answer
2012-09-07T09:25:48-0400
V=4.18*10^20L
p=1.56 g/ml = 1560kg/L
m=pV

m=4.18*10^20L
* 1560kg/L=6.5208*10^23
m of NaCl= 6.5208*10^23 * 0.043 =2.804*10^22 kg